Mary has $2.15 in change consisting of 11 more dimes than nickels. Find the number of nickels and dimes. how do you set up this problem and solve it?
Answer by jorel1380(3719)

Assuming Mary has only nickels and dimes,then:
nickels=x
dimes=x+11
5x+10(x+11)=215
15x+110=215
15x=105
x=7
x+11=18
Mary has 7 nickels, and 18 dimes..